OpenStudy (aravindg):

The greatest acceleration ordeceleration that a train may have is a.find minimum time in which the train may reach from one station to other separated by a distance d

OpenStudy (jamesj):

So what you want to do obviously is accelerate to the midpoint and decelerate from the mid-point to the end. Given that, how long does it take to reach the mid-point?

OpenStudy (aravindg):

?

OpenStudy (jamesj):

How far is the mid point from the first station? d/2. How far does the train travel if it starts at rest and accelerates at a constant rate of a?

OpenStudy (jamesj):

c'mon Aravind. How far does a body move if it starts at rest and accelerates at a constant rate of a for a time of t?

OpenStudy (jamesj):

If you're not going to stay and engage with me I don't know why I bother to try and help you. This isn't the first time you walk away mid-conversation and it gets pretty boring for me if you're not going to make an effort. Sorry.

OpenStudy (aravindg):

srry james i will try not to repeat it

OpenStudy (jamesj):

ok. So where did you get with this question?

OpenStudy (jamesj):

Aravind, talk to me here, not in other people's questions.

OpenStudy (aravindg):

k

OpenStudy (jamesj):

Have you finished this question? Where are you with it?

OpenStudy (aravindg):

i am doing another currently

OpenStudy (aravindg):

i am discussing with turing can u also cm

OpenStudy (aravindg):

http://www.twiddla.com/712316

OpenStudy (jamesj):

well, I'd like to see you finish one question (with my help) before I engage with you on another. I'm unhappy that I've never seen you reach the end of a problem. When you're ready, let's finish this problem. Then I will help you with a new problem. But for now I'm not interested in helping you with a new one. When you're ready, let me know here.

OpenStudy (aravindg):

K

OpenStudy (aravindg):

for minimum time acceleration is a

OpenStudy (aravindg):

james?

OpenStudy (jamesj):

Oh, ok, you want to do this now. So the train accelerates to the half way point d/2, then decelerates.

OpenStudy (jamesj):

The first question is, how long does it take to reach d/2 if it starts from rest. So if it accelerates at a constant rate, then velocity v at time t is v = at. What then is the displacement at time t?

OpenStudy (jamesj):

If s(t) is the displacement, we want to know the time t such that s(t) = d/2. Hence we need a formula for s(t). What is that formula, given that v(t) = at ?

OpenStudy (jamesj):

And you're not here? Where are you?

OpenStudy (jamesj):

For the record \[ s(t) = s(0) + \int_0^t v(u) \ du \] \[ = s(0) + \int_0^t au \ du \] \[ = s(0) + \frac{1}{2} at^2 \] As s(0) = 0, hence \[ s(t) = \frac{1}{2}at^2 \]

OpenStudy (aravindg):

ya

OpenStudy (aravindg):

so 1/2 at^2 is the answer??

OpenStudy (jamesj):

s(t) = 1/2 . at^2. So how long does it take the train to reach the mid-point d/2 ?

OpenStudy (aravindg):

at^2

OpenStudy (jamesj):

No. At the mid-point s(t) = d/2, as s(t) is the expression for displacement at time t. So the question is: for what value of t is this the case?

OpenStudy (aravindg):